The premium on a June 17 British pound call option with a strike price of $1.2560 when the spot rate is $1.2620 is quoted as $0.02. The time value of this option is.

The premium on a June 17 British pound call option with a strike price of $1.2750 when the spot rate is $1.2620 is quoted as $0.025. The intrinsic value of this option is.

Your firm has an accounts receivable worth C$200,000 due in six months. The firm buys 2 June 16 Canadian dollar put options with a strike price of $0.9750 at a premium of $0.0075 to hedge against the exchange risk. If the spot rate in June settles above the strike and current interest rate dollar 2%, the outcome of the hedge will be exactly $193,483. Yes or no? explain
Answer #1

1.
=0.02-MAX(1.2620-1.2560,0)=0.014

2.
=MAX(1.2620-1.2750,0)=0

3.
No
Profit from put=200000*MAX(0.9750-0.9750*1.02,0)-200000*0.0075=-1500
The outcome of the hedge=200000-1500=198500